OpenStudy (anonymous):

As cars passed a checkpoint, the following speeds were clocked and recorded. Speed (mph): 55 62 61 54 68 72 59 61 70 15. Find the mean and the range of the data set. 16. What are the minimum, first quartile, median, third quartile, and maximum of the data set?

OpenStudy (anonymous):

The mean is the average of all the numbers in a dataset. To get it, you add all the numbers in the set together, and then divide it by the number of numbers in the set. (55+62+61+54+68+72+59+61+70)/9 The range of a set of data is the difference between the largest and smallest values. 72-54 The minimum value of a data set is the least value in the set. 54 The maximum value of a data set is the greatest value in the set. 72 1st Quartile is the median of the lower half of the data, the 3rd quartile is the median of the upper half of the data. The interquartile range of a data set is the distance between the two quartiles. Interquartile range = Q3 – Q1.

OpenStudy (anonymous):

54,55,59,61,61,62,68,70,73 lowerhalf is 54,55,59,61; upperhalf is 62,68,70,73 Q1=(55+59)/2 Q3=(68+70)/2

OpenStudy (anonymous):

Thank you for your help.

OpenStudy (anonymous):

The interquartile range is 13. is that right?

OpenStudy (anonymous):

Sorry, I midread the dataset, it should have been 54,55,59,61,61,62,68,70,72; the interquartile range should be 12, as Q3=69, Q1=57 and 69-57=12.